OpenStudy (anonymous):

The perimeter of a rectangle is 72 centimeters. The base is 2 times the height. What are the dimensions of the rectangle

OpenStudy (anonymous):

2(b+3b)=72 2b+6b=72 8b=72 base=9 cm. and height=27 cm. so area = 243 sq. cm.

OpenStudy (texaschic101):

formula for perimeter of rectangle : P = 2L + 2W or in this case P = 2B + 2H P = 72 B = 2H now we sub 72 = 2(2H) + 2H 72 = 4H + 2H 72 = 6H 72/6 = H 12 = H now sub 12 in for H B = 2H B = 2(12) B = 24 check.. P = 2B + 2H 72 = 2(24) + 2(12) 72 = 48 + 24 72 = 72 (correct) Height (H) = 12 cm Base (B) = 24 cm Area = BH Area = 24 * 12 Area = 288 cm^2